OpenStudy (anonymous):

yes. still doing this problem. 10 gallons of diesel fuel C16H34 density 0.7g/ml C16H34+O2 -> CO2+H2O How many CO2 kg is produced? I just need help on finding how many CO2 is produced..

OpenStudy (anonymous):

Im at the point where I found the Molar Mass, Ratio, Balanced the equation, I just am confused on how to exactly set up the rest from there.

OpenStudy (australopithecus):

Can you provide me with the molecular mass, and the balanced equation and moles of C16H34 in 10 gallons ?

OpenStudy (anonymous):

I converted the 10 gal to ml which is : 37854.1 ml or 3.78 x 10^4 Then I have the molar mass of C16H34 which is : 226.446 g I balanced the equation to : 2 C16 H34 + 49 O2-----> 32 CO2 + 34 H20 The Molar ratio is: 2 to 29 to 32 to 34 Now what!!! :)

OpenStudy (australopithecus):

Memorize this formula \[Moles = \frac{grams}{Molecular Mass}\] NOTE molar mass has units g/mol 37854.1 mL * 0.7g/mL = 26500g notice that mL cancels out and you are left with grams now convert to moles using the formula I provided you with that you should memorize 26500g/226.446 g/mol = then divide by 2 and multiply by 32 to convert to moles of CO2

OpenStudy (australopithecus):

then just use the formula again to convert to grams of CO2, then divide by 1000 to convert grams of CO2 to kg of CO2

OpenStudy (anonymous):

okay so I got 1872.41 after dividing by 2 and multiplying by 32 now ill convert to grams of CO2...

OpenStudy (anonymous):

because the 1872.41 is moles..

OpenStudy (australopithecus):

yes notice that grams cancel out and you are left with moles as a unit

OpenStudy (anonymous):

so i got 0.008268 kg CO2

OpenStudy (australopithecus):

44.0095g/mol is the molecular mass of CO2 so use the formula 1872.41mol = x/44.0095g/mol basic algebra you want to isolate x so multiple both sides by 44.0095g/mol x = 1872.41mol*44.0095g/mol notice that moles cancel out (look at units like numbers when you divide them they just become 1 so they disapeer) x = 82404g/1000 = 8.2kg

OpenStudy (anonymous):

oh okay I just Put the molar mass of the other one instead of CO2 …got it now.

OpenStudy (australopithecus):

well I hope you understand how I got the answer, this stuff is pretty fundamental. If you have anyother questions about how I came to the solution ask.

OpenStudy (australopithecus):

just make sure units cancel and you shouldn't have any problems

OpenStudy (anonymous):

Yeah I totally understand canceling items but Its hard to figure what to do next sometimes.. But ill work it through once more from the beginning so thank you

OpenStudy (australopithecus):

yeah its not a problem it is really systematic once you do it a couple times you wont have a problem with it